On this page, we list the major changes that affected the instrument status during the lifetime of SINFONI. Note that these are in general changes that also influence the calibration database to be used.

  • 30 June 2019: SINFONI has been decommissioned at the end of June 2019. The last science observations were taken on June 25. The SINFONI entry has been removed from Phase 1.
  • 21 April 2019: The SINFONI intervention was successful and the MACAO is repaired with a new membrane. SINFONI is now fully operational.
  • 10 April 2019: Cooling water was detected inside MACAO that came from a leak in the cooling lines of the Cassegrain TCCD. Because of this leaking problem, SINFONI has been dismounted from UT3. Service and visitor observations are therefore suspended until MACAO is fixed.
  • 6 Febuary 2019: The SINFONI transfer was successful, and the recommissioning was carried out in the first two weeks of February. After moving to UT3, the system performance is very similar to that at UT4. Some changes include: 1). the slit positions are shifted by 4 pixels in the positive x-direction; 2). the detector blind spot is  shifted by 4 pixels more into slit #16, which is similar to the previous position before the 2018 intervention.
  • 19 Febuary 2018: The SINFONI intervention was successful, and the recommissioning was carried out in the beginning of February. The damaged preoptics motor has been replaced and the new motor is functioning well. Maintenance tasks were also performed on many parts of SINFON, such as the MACAO AO system. After the intervention, the system performance is, overall, similar to that before the intervention. Some changes include: 
  1. Installation of a new membrane mirror. Slight improvement in internal focus in shorter wavelengths (J and H) with no impact in K.
  2. The orientation of SINFONI with respect to the sky has been recalculated and numerically adjusted by 0.26 degrees for the NGS mode and 0.5 degrees for the LGS mode.
  3. The locations of the spectra are shifted slightly on the detector, by about 4 pixels in x and 1-2 pixels in y.
  4. The detector blind spot is now shifted 4 pixels farther away from the center of the field of view, while shifting more into slit #1 (bottom slit in the FOV).
  5.  The pipleline reduces the new data without issues, except for the distortion recipe for the J-band NorthSouth data, where the bp_dist-min_cut parameter should be set to 0.05 (instead of 0.1). This value is found to work for data before and after the intervention, and will be implemented in the next release of the pipeline.
  • 04 November 2017: Recently UT4 has been experiencing intermittent issues with image elongation, affecting the data taken by SINFONI. While the cause of this occasional elongation is under investigation.
  • 30 June 2017: The laser is now back in operation with SINFONI.
  • 18 June 2017: Due to a technical problem, SINFONI observations with the laser are be available until at least the end of June.
  • 4 December 2016: SINFONI underwent successfully recommissioning following the installation of the deformable secondary mirror on UT4.
  • 23 September 2016: SINFONI will be offline from approx 9 October onwards until the end of November due to M2 installation.
  • 13 September 2016: The MACAO shutter for the AO unit was replaced successfully in a mini intervention on 12/9.
  • 14 March 2016: As SINFONI underwent upgrade. In a nutshell, the performance improvements include:
  1. A significant improvement in the overall J-band throughput
  2. Increased spectral resolution in all bands
  3. A spot of dead pixels which previously affected the centre of the field of view are now shifted mostly out of the way
  4. The vignetting of slit 32 at the top of the field of view is now much reduced
  5. An improved AO performance of ~10% in Strehl
  6. No significant negative impact on the system performance.
  • 2 June 2014: grating wheel speeds have been lowered: the normal speed is 14 encoders/ms, and the fine speed is 5 enc/ms. A grating change between the 2 most separated gratings (J and H+K) will now take 3.7 minutes.
  • 21 September, 2013: Post recommissioning, all filter/grating combinations and observing modes are available. Calibration database is being repopulated. Dark exposures and linearity tests are resumed.
  • 2 September, 2013: A new Pupil-Tracking mode will be offered in Period 93, enabling the Spectrocopic Angular Diffrential Imaging (S-ADI) technique to be used for certain targets. Please see the SINFONI Manual for more details.
  • 17 September, 2013: SINFONI intervention was successful. Four motors were replaced (grating, filter, pre-optics and sky-spider). The coating of the 25mas pre-optics lens was discovered to be degraded (likely during warm up) and this was successfully polished off. The detector is displaced 5 pixels compared to before (slit positions are shifted 5 pixels in the positive x-direction.)
  • 24 July, 2013: Filter wheel is still stuck in K. Due to the delicate state of the filter wheel and the substantial risk of getting it stuck while moving it (thus rendering the instrument inoperable), the Instrument Operation Team has decided to keep SINFONI in K until 12 August,
  • 10 June, 2013: Filter wheel is still stuck in K, and the aforementioned shift in wavelength is still happening.
  • 12 May, 2013: The grating has displaced slightly, resulting in a shift of 2-3 pixels in the dispersion direction compared to normal. Since the science and the arcs are taken with the same shift, the data can be calibrated.
  • 15 April, 2013: We are experiencing problems with filter wheel. Currently the only solution is to keep the filter in a fixed position under simulation mode. As a result darks exposures cannot be taken. Service observations are therefore suspended until this problem is fixed.
  • 1 March, 2013: The new laser PARLA has been successfully commissioned and is now in operation with SINFONI, with LGS available on most nights. The on-sky power is about 6W.
  • 26 Feb, 2013: The strong noise pattern as reported in November 2012 is still occassionally seen. The pattern is not fixed in position, but moves around the detector, with a RMS > 10 ADUs. Observers should look out for this.
  • 20 Feb, 2013: The motors for the Grating and Filter wheels are failing. Intervention is expected in September 2013. In the meantime, in order to prolong the lives of the motors and minimise further damage, the speed of the motors have been drastically decreased, such that a grating/filter change will now take about 10 minutes.
  • 29 November, 2012: Following troubleshooting by Instrumentation, the strong noise pattern has mainly gone. Service mode OBs with faint targets are put back into the queue.
  • 11 November, 2012: Noise pattern is seen in science and calibration data. This is being investigated. Until this problem is solved, for service mode observations preference will be given to bright targets.
  • December 2010: With the start of P87 Phase 2, the new template 'SINFONI_ifs_cal_GenericOffset' has been made available, to allow efficient PSF observations in LGS-noTTS mode (Seeing Enhancer mode); via this template, PSF calibrations data can be taken within the same OB as the science data.
  • May 2010: For LGS mode operations, SINFONI has received a new STRAP head that should help to overcome the problem with the central obscuration of the old STRAP head experienced under very good seeing conditions (i.e., due to its narrow PSF, the TTS was lost in the central obscuration).
  • April 2005: The data taken with SINFONI are now visible from the ESO Archive.
  • July 2004: SINFONI saw first light at UT4.